OpenStudy (anonymous):

What is the solution of the system? Use elimination. 3x – 4y = –24 x + y = –1 A. (–4, 3) B. (0, 6) C. (3, 4) D. (4, 3)

OpenStudy (anonymous):

3x – 4y = –24 (1) x + y = –1 (2) ->x=-1-y (3) let (3) into (1) so 3*(-1-y)-4y=-24 ->y=3 ->x=-1-3=-4

Parth (parthkohli):

\(\Large \color{MidnightBlue}{\Rightarrow 4x + 4y = -4 }\) Add both equations. \(\Large \color{MidnightBlue}{\Rightarrow4x = -28 }\) \(\Large \color{MidnightBlue}{\Rightarrow x = -4 }\) Lol we know by now that x = -4 so it is A.

OpenStudy (amorfide):

try to get the same amount of x's in each equation or same amount of y's to do this you have to multiply one equation so lets say i want the same x's multiply equation 2 by 3 now i have two equations like this 3x – 4y = –24 3x + 3y = –3 now you will subtract equation 2 from 1

Parth (parthkohli):

By now, all the options but a are eliminated.

OpenStudy (anonymous):

no! it's A